OpenStudy (anonymous):

If all numbers are defined to be positive integers of the form 4n + 1 (i.e. 1,5,9,13,17, etc), find the first six primes. I think they're 5, 9, 13, 17, 21, and 29, but that seems wrong for some reason.

OpenStudy (anonymous):

Yeah, but it's excluded in this case.

OpenStudy (swissgirl):

9 aint a prime number

OpenStudy (swissgirl):

9=3*3 21=7*3

OpenStudy (anonymous):

There is no 3 in this set of integers. ;)

OpenStudy (swissgirl):

ohhhhhhh i get ur question

OpenStudy (anonymous):

Seems obvious that those are the first six primes, but my intuition is warning me that something is awry.

OpenStudy (swissgirl):

@dan815 what do you think?

OpenStudy (dan815):

yeah

OpenStudy (dan815):

numbers of 4n+1 numbers 1,5,9,13,17,21,25 5 is a prime because onl 1 and 5 go into it same wth 9 since 1 5 doent go into it 13 since 5 and 9 dont... and so on

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

They need an openstudy area for programming. :)

OpenStudy (swissgirl):

There is ... On the upper left hand side next to the amount of fans it says find more subjects ... Check there